OpenStudy (anonymous):

solve using substitution method: -x-3y=-20, 5x+3y=40

OpenStudy (lgbasallote):

go guyss! :DD

OpenStudy (anonymous):

\[5 (20-3 y)+3 y=40 \]\[y=5 \]

OpenStudy (anonymous):

-x-3y=-20 -x=-20+3y x=20-3y 5(20-3y)+3y=40 100-15y+3y=40 100-12y=40 -12y=-60 y=5

OpenStudy (anonymous):

Then plug the y value back into the one of the equations at the top, and you should get the answer for x

OpenStudy (anonymous):

From first equation x=20-3y. Now substitute this in second one for x. 5(20-3y)+3y=40 Gives 100-12y=40 Implies 12 y=60 which means y=5 Now we have x=20-3y=20-3(5)=5 So answer is x=5 and y=5

OpenStudy (anonymous):

Thank you so much. I have watched so many videos trying to understand this and seeing both of you explaining it makes so much sense now. :)